OpenStudy (anonymous):

Demonstrate the proof of a polynomial identity through an algebraic proof and a numerical proof. (Some Examples) (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d) I am unclear on what I am supposed to do.

OpenStudy (sleepyjess):

I believe you have to create your own proof. I did something like this once

OpenStudy (anonymous):

By just plugging number in or is their a better way than plug 'n chug?

OpenStudy (anonymous):

*numbers

OpenStudy (sleepyjess):

Well, basically I just took one of those, like (x+y), then cubed it, \((x+y)^3\), found what that equaled, then plugged in a number for x and y, then confirmed the identity

OpenStudy (anonymous):

So say \[(x+y)^3\] where x=3 and y=6 I would add x+y and then cube it correct? So \[9^3\] Which = 729

OpenStudy (sleepyjess):

Well, actually,\((x+y)^3 = (x+y)(x+y)(x+y)\). But I believe that would come out to be the same once you substitute, but first we have to find what \((x+y)^3\) is in terms of x and y

OpenStudy (anonymous):

What?

OpenStudy (sleepyjess):

(x+y)^3 = (x+y)(x+y)(x+y)

OpenStudy (anonymous):

To make it a smaller number I'm going to use \[(x-y)^3\] So \[(6-3)^3=(6-3)(6-3)(6-3)\] \[6-3=3\] Soo \[(3)(3)(3)=27\]

OpenStudy (sleepyjess):

Do you know FOIL

OpenStudy (anonymous):

First Outside Inside Last

OpenStudy (anonymous):

Ahhh your joking

OpenStudy (sleepyjess):

Yes, we need to use that with \((x+y)(x+y)(x+y)\) without substituting in any numbers

OpenStudy (anonymous):

You know sometimes I despise FOIL

OpenStudy (anonymous):

\[x^2-xy-xy+y^2(x-y)\]

OpenStudy (anonymous):

Then rinse and repeat

OpenStudy (anonymous):

\[x^3-x^2y-x^2y+xy^2\]

OpenStudy (sleepyjess):

Why -xy-xy?

OpenStudy (anonymous):

\[(x-y)(x-y)\] First:x*x=x^2 Outside:x*-y=-xy

OpenStudy (anonymous):

Inside and last are the same thing

OpenStudy (sleepyjess):

Oh, I thought we were doing x+y :P

OpenStudy (anonymous):

Haha Yea sorry bout didn't like the 729, 27 seemed more amiable

OpenStudy (anonymous):

So then \[x^3−x^2y−x^2y+xy^2\] is correct?

OpenStudy (anonymous):

From here I plug in 6 for x and 3 for y yes?

OpenStudy (sleepyjess):

Remember, we need to multiply each term by x and -y \(x^2 * x\\-xy*x\\-xy*x\\+y^2 * x\\+x^2*y\\-xy*-y\\-xy*-y\\+y^2*-y\)

OpenStudy (anonymous):

\[3^2*3\] ?

OpenStudy (sleepyjess):

\(\color{blue}{\text{Originally Posted by}}\) @Prometheus777 So then \[x^3−x^2y−x^2y+xy^2\] is correct? \(\color{blue}{\text{End of Quote}}\) All you did here was multiply it all by x

OpenStudy (anonymous):

What I did was FOIL \[(x-y)^3\] which gave me \[(x-y)(x-y)(x-y)\] from there I foiled \[(x-y)(x-y)\] which gave me \[x^2-xy-xy+y^2\] So far so good?

OpenStudy (sleepyjess):

Yes, that is correct so far

OpenStudy (anonymous):

Alright, with \[x^2-xy-xy+y^2\] I proceeded to FOIL that against \[(x-y)\] \[x^2-xy-xy+y^2*(x-y)\] Which gave me \[x^3-x^2y-x^2y+xy^3\]

OpenStudy (anonymous):

So do I plug in the numbers or is there another step?

OpenStudy (anonymous):

When I plug 6 in for x and 3 in for y I get 162...

OpenStudy (sleepyjess):

I got \(x^3+2xy^2-y^3\)

OpenStudy (anonymous):

How the what?

OpenStudy (anonymous):

I understand all but the last part the \[-y^3\]

OpenStudy (sleepyjess):

\(x^2 * x=x^3\\-xy*x=-x^2y\\-xy*x=-x^2y\\+y^2 * x=xy^2\\+x^2*y=+x^2y\\-xy*-y=xy^2\\-xy*-y=xy^2\\+y^2*-y=-y^3\) \(x^3-x^2y-x^2y+xy^2+x^2y+xy^2+xy^2-y^3\) combine like terms \(x^3-3x^2y+3xy^2-y^3\) I feel like I did something wrong now...

TheSmartOne (thesmartone):

You did it correct Jess :)

OpenStudy (anonymous):

I plug in 6 for x and 3 for y and it works :-) Now can you explain how you got that?

TheSmartOne (thesmartone):

You could even use pascal's triangle: